nekoguy:

A right triangle has a leg of 13 cm and a hypotenuse of 21 cm. What is the length of the other leg? Round to the nearest tenth. 8.0 cm 16.5 cm 24.7 cm 272.0 cm pleeeeease help me right "meow"

darkknight:

since this is a right triangle you can use the pythagoreon theorem (a^2+b^2= c^2) c is the hypo and a and b are legs

kittybasil:

As per the Pythagorean Theorem mentioned previous:

  • "A right triangle has a leg of 13 cm" this can be \(a\) or \(b\) since both represent leg values. I'm going to use \(b\), so \(\color{lightcoral}{b=13\text{ cm}}\)
  • "a hypotenuse of 21 cm" which is referred to by \(c\), so \(\color{plum}{c=21\text{ cm}}\)
Now we construct the expression generated by the formula:\[a^2+\color{lightcoral}{b}^2=\color{plum}{c}^2\text{ becomes the following: }a^2+(\color{lightcoral}{13})^2=(\color{plum}{21})^2\]Now let's isolate the variable before we simplify anything.\[a^2\cancel{+13^2\color{lightgreen}{-13^2}}=21^2\color{lightgreen}{-13^2}\]\[a^2=21^2-13^2\]Simplifying, we square these values and get:\[a^2=441-169=272\]Now we solve for \(a\) by isolating it. By this, we do:\[\sqrt{a^2}=\sqrt{272}\]\[a=\sqrt{272}\]What answer did you get?

TheSmartOne:

Alright folks. Enough is enough. This post has been spammed as it is and this conversation is not going anywhere and neither is it very beneficial in that it's providing clarity to any of you. There are multiple issues at hand and it's going to take a while to address them all. So, we're going to ask you all to go take a break outside. We can see deleted posts, we (the mods) will respond shortly. In the meanwhile, STOP RESPONDING. If you want a forced break (aka suspension) then continue to post away. No more responses from here on. Just wait for us to respond (and that'll take a while).